Fermi National Laboratory
 

Accelerator Update for December 6, 2006 - December 8, 2006

 

Wednesday December 6

The day shift began with Operations monitoring store 5108 (106.27E30) and a Recycler stash (6.1E10), with the Antiproton Source (Pbar) stacking (116.3E10), with Linac, Booster, the Recycler, and the Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

The Recycler completed a stash of antiprotons at 9:16 AM. The first transfer of antiprotons did not make it into the Recycler. Also, a MI RF station (MIRF12) tripped off during the transfers. This caused the NuMI beam permit to drop. It reset okay.

At 9:24 AM, Operations reported that when Pbar resumed stacking after the stash, operators noticed that a Septa (D:ISEP) wasn’t pulsing. Operators contacted an expert who investigated the problem, but the Septa resumed pulsing before he could find the cause.

The Accelerator Division Head gave permission to run beam to the Switchyard (SY) 120 Absorber at 10:07 AM. Operations still needs a set of Run Conditions before running any beam.

At 10:59 AM, store 5108 was lost due to a power supply trip (T:SD). There was no quench.

At 11:06 AM, a fire alarm posted for the TeV sector C4 area. Coincidentally, the T:SD power supply was located at C4.

Cryo system technicians took advantage of the downtime at 11:18 AM and began a valve adjustment on the TeV sector A3 wet engine.

At 11:36 AM, operators discovered charred components inside the T:SD regulator. Power supply experts began their repairs.

The left and middle pictures show the regulator. The picture on the right shows the transductor electronic chassis, which was above the regulator.

Operations had CDF and D0 prepared for access by 11:55 AM.

Power supply experts had all the T:SD equipment replaced by 12:09 PM.

The Recycler completed a stash at 1:14 PM.

Operations reported at 1:58 PM that they had received SY 120 Run Conditions.

SY experts began starting up the beam line around 2:13 PM.

D0 came out of access at 3:27 PM.

Operations began putting the TeV through an extended dry squeeze at 3:30 PM.

At 4:04 PM, Operations reported that T:SD needed more work.

At 4:24 PM, Operations reported a Pbar link problem. Operators investigated and found a bad 5-volt power supply at AP-10.

The Recycler completed a stash at 5:24 PM. The first transfer of antiprotons did not make it out of the Accumulator. At the end of the stash, D:ISEPV did not pulse again. Operators caught the problem, cycled the power, and stacking resumed.

Operations put the TeV through another dry squeeze at 5:28 PM.

Operations reported at 6:14 PM that all Meson Primary safety system tests were successfully completed.

Operations reported at 6:21 PM that the MI slow spill extraction Septa was operating. (This is how beam gets sent to SY.)

Operations loaded a 14x0 wet squeeze into the TeV at 7:02 PM (14 bunches of protons and zero bunches of antiprotons).

Operations put the TeV into shot setup at 7:37 PM.

At 7:55 PM, a Booster RF station (BRF6) tripped off on a ground fault indication. An expert investigated and at 8:45 PM said the ground fault was real.

Operations established store 5111 at 9:47 PM with an initial luminosity of 135.78E30.

Thursday December 7

The midnight shift began with Operations monitoring store 5111 (87.50E30) and a stash (51.5E10), with Pbar stacking (53.7E10), and with NuMI and MiniBooNE taking beam.

At 2:01 AM, the I- Source tripped off on an overcurrent indication. It reset okay.

At 5:39 AM, the I- Source went down again and this time it would not reset. Operations called in experts to help switch over the H- Source. Until the H- Source comes on line, there is no beam to downstream machines.

At 6:51 AM, Operations reported that the H- Source wasn’t warm. It will take several hours before it can deliver any beam.

Operations prepared the Booster for access at 8:18 AM.

The Recycler completed a stash at 9:47 AM.

The Neutron Therapy Facility (NTF) reported at 9:55 AM that they hoped to begin treating patients at 11 AM.

At 10:05 AM, Operations received a report of a LCW leak in MI-10 and the smell of burnt electronics. Operators locked out the power supplies and valved off the LCW.

MiniBooNE reported at 10:10 AM that they were switching to a horn off run mode.

A Preaccelerator expert began tuning the H- Source at 10:12 AM.

Booster came out of access and was turned on at 12:15 PM.

MI came back up (minus MI-10).

Pbar resumed stacking at 12:23 PM.

NuMI and MiniBooNE Resumed taking beam by 12:48 PM.

NTF reported at 2:45 PM that they had finished treating their last patient for the day.


The first pulse of slow spill to SY occurred at 3:30 PM.

The Recycler completed a stash at 5:46 PM.

Operations reported at 5:46 PM that NuMI beam was again falling out of MI during the stash and after the stash.

Operations terminated store 5111 at 8 PM.

Operations prepared CDF and D0 for an access at 8:18 PM.

At 8:22 PM, NuMI beam was falling out of MI during the stash.

The Recycler completed a stash at 8:31 PM.

CDF came out of access at 10:16 PM.

At 11:08 PM, NuMI beam was falling out of MI during and after the stash.

The Recycler completed a stash at 11:18 PM. A MI expert contacted Operations and suggested a fix. Operations loaded a one-shot at 11:38 PM from the Accumulator to Recycler aggregate and this fixed the problem.

Friday December 8

The midnight shift began with Operations monitoring the stash (246.9E10), with Pbar stacking (28.1E10), with D0 in access, and with NuMI and MiniBooNE taking beam.

Operations set up the Recycler to stash at 1:48 AM.

At 2:18 AM, Operations reported that NuMI was falling out of MI during the stash and continued after the stash had ended. Operators tried the trick the last shift had used, but it didn’t work initially, but then it did. However, when operations raised the intensity, NuMI beam again fell out of MI. Operators tuned and were able to maintain the higher intensity with no fall out.

D0 came out of access 3:40 AM.

Operations put the TeV through a dry squeeze at 4:42 AM.

Operations put the TeV into shot setup at 5:15 AM.

Interlock technicians reported at 6:45 AM that they had completed the Meson safety system tests.

Operations established store 5113 at 6:59 AM with an initial luminosity of 177.58E30.

Operations prepared MI for an access at 7 AM. Instrumentation experts will work on the Recycler DCCT.

At 7:09 AM, Operations reported that NuMI beam was falling out of MI during the shot setup. The Crew Chief made an adjustment to the low level RF and this stopped the fall out. He set everything back to normal after the shots. He didn’t think what he did was a real fix.

The Plans for Friday

The plans for today are to stack, stash, and store. The Run Coordinator will allow a few studies during the day. MI experts will work on the NuMI fallout problem.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/11/2006    email Fermilab